OpenStudy (anonymous):

What are 2 ways to write the number as a sum of fractions with the same denominator 1 1/4

OpenStudy (zzr0ck3r):

\(1\frac{1}{4}=1+\frac{1}{4}=\frac{1}{1}+\frac{1}{4}=\frac{4}{4}+\frac{1}{4}=\frac{4+1}{4}=\frac{5}{4}\)

OpenStudy (anonymous):

Thanks